A sphere of radius 2 ft is dropped into a rectangular tank that is 10 ft long and 8 ft wide. To the nearest tenth, how much does this raised the level of water in the tank?
Answer by checkley77(12844)

V=4/3PIR^3
V=4/3*3.14*2^3
V=4/3*3.14*8
V=4/3*25.12
V=100.48/3
V=33.49 FT^3
10*8=80 FT^2 IS THE SURFACE AREA OF THE TANK.
33.49/80=.42 FT. OR 5 INCHES IS THE RISE IN WATER LEVEL (ASSUMING THE SPHERE SINKS BELOW THE SURFACE..